# A Multi-fidelity Estimator of the Expected Information Gain for Bayesian Optimal Experimental Design

### v1

@@ 6 Conclusions
This paper introduced MF-EIG, a novel EIG estimator for OED problems that accommodates general multi-fidelity ensembles of models under the ACV framework. Notably, this estimator is unbiased with respect to the high-fidelity mean, and minimizes variance under a given computational budget. We achieved this by first presenting a reparametrization of the EIG that formulates its expectations to be independent of the data models, a requirement for compatibility with the ACV framework. This reparameterization also provided insights into the inner mechanics of EIG estimation and the notion of data model fidelity. We then provided specific examples of NMC-style utility models under additive independent and scaled data noise forms, and proposed practical enhancements via inner-loop sample size optimization and sample reuse techniques. We demonstrated the MF-EIG estimator in two numerical examples: a one-dimensional nonlinear benchmark where we made extensive comparisons under different settings, and a multi-dimensional turbulent flow problem involving the calibration of SST turbulence closure model parameters within the RANS model where we also provided explanations to the behavior of results through physical principles and insights. We validated the estimator’s unbiasedness and variance reduction through these examples, where MF-EIG achieved one- to two-orders-of-magnitude variance reductions, with moderate additional improvements from optimizing the inner-loop samples and substantial improvement from reusing the inner-loop samples. We implemented MF-EIG using the MXMCPy toolbox, and made our code available at: https://github.com/tcoonsUM/mf-eig/.

### v2

@@ 6 Conclusions
This paper introduced MF-EIG, a novel EIG estimator for OED problems that accommodates general multi-fidelity ensembles of models under the ACV framework. Notably, this estimator is unbiased with respect to the high-fidelity mean, and minimizes variance under a given computational budget. We achieved this by first presenting a reparametrization of the EIG that formulates its expectations to be independent of the data models, a requirement for compatibility with the ACV framework. This reparameterization also provided insights into the inner mechanics of EIG estimation and the notion of data model fidelity. We then provided specific examples of NMC-style utility models under additive independent and scaled data noise forms, and proposed practical enhancements via inner-loop sample size optimization and sample reuse techniques. We demonstrated the MF-EIG estimator in two numerical examples: a one-dimensional nonlinear benchmark where we made extensive comparisons under different settings, and a multi-dimensional turbulent flow problem involving the calibration of SST turbulence closure model parameters within the RANS model where we also provided explanations to the behavior of results through physical principles and insights. We validated the estimator’s unbiasedness and variance reduction through these examples, where MF-EIG achieved one- to two-orders-of-magnitude variance reductions, with moderate additional improvements from optimizing the inner-loop samples and substantial improvement from reusing the inner-loop samples. Based on these results, we strongly recommend utilizing inner-loop sample reuse. Where practical, we also recommend optimizing for low-fidelity inner-loop samples sizes, though its impact appears to be less pronounced than that of sample reuse. We implemented MF-EIG using the MXMCPy toolbox, and made our code available at: https://github.com/tcoonsUM/mf-eig/.